Brass is an alloy of copper and zinc. Varying the mass percentages of the two metals produces brasses with different properties. A brass called yellow zinc has high ductility and strength and is 34.0%−37.0% zinc by mass.

(a) Find the mass range (in g) of copper in 487 g of yellow zinc.

(b) What is the mass range (in g) of zinc in a sample of yellow zinc that contains 48.5 g of
copper?
Answer #1

(a) Mass range of Zinc in yellow bronze = 34 - 37%

Mass range of Cu = 100% - mass % of Zinc = 66% - 63%

Mass range of Cu = 0.66 * 487 = 321.42 g, 0.63 * 487 = 306.81 g

(b) Assuming 66% Zn and 34% Cu combination:

Mass of Zn = (66 / 34) * 48.5 = 94.15 g

Assuming 63% Zn and 37% Cu combination:

Mass of Zn = (63 / 37) * 48.5 = 82.58 g

Mass range = 82.58 - 94.15 g